A március 17.-i valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Balázs István  9  9  6  9  6  5  44
Dékány Tamás  9  7  6  8  9  5  44
Dudás János  2  9  0  9  0  5  25
Gyenizse Gergő  9  9  9  9  5  5  46
Gyöngyösi Zsolt  9  9  9  9  9  0  45
Szakács Nóra  9  9  5  8  6  0  37
Szalánczi Szabolcs  0  7  0  0  0  0  7